OpenStudy (anonymous):

(2x-4)(2x-1)= (1 point)

OpenStudy (anonymous):

FOIL.

OpenStudy (anonymous):

4x^2-9x+4

OpenStudy (anonymous):

2,and \[\frac{ 1 }{ 2}\]

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Are you trying to solve for x?

OpenStudy (anonymous):

I have to solve the equation using the zero product property.

OpenStudy (anonymous):

Ohh ok. SO set both pieces equal to zero. Then solve for x.

OpenStudy (anonymous):

\[(2x-4)(2x-1)=0\]

OpenStudy (anonymous):

\[2x-4=0 \] and \[2x-1=0\]